# Prompt-CAM: A Simpler Interpretable Transformer for Fine-Grained Analysis

### v1

@@ 4.3 Further Analysis and Discussion
Prompt-CAM on different backbones. Figure 7 illustrates that Prompt-CAM is compatible with different ViT backbones. We show the top three attention maps generated by Prompt-CAM using different ViT backbones on an image of “Scott Oriole.” These maps highlight consistent identification of traits for species recognition, irrespective of the backbones. Please see the caption and Suppl. for details.
@@
Prompt-CAM can detect biologically meaningful traits. Figure 1 illustrates that Prompt-CAM can consistently identify traits from images of the same species to recognize it (e.g., breast color and wing pattern in the first and second columns of Figure 1.b for Baltimore Oriole). Another species like Bronzed Cowbird may see some of its traits from the images (e.g., head color and wing pattern in Figure 1.c) but not all of them (e.g., breast color), thus is unable to claim the images as Bronzed Cowbird. These results demonstrate Prompt-CAM’s ability to recognize species in a biologically meaningful way.
Figure 8: Trait manipulation. The top row shows attention maps for a correctly classified Red-winged Blackbird image. In the second row, the red spot on the bird’s wings was removed, and Prompt-CAM subsequently classified it as a Bronzed Cowbird, as depicted in the reference column.
Figure 9: Focused Prompt-CAM (subsection 3.4) applied to pair-wise comparison. Given different references, Prompt-CAM looks at different (numbers of) traits to recognize the input image. For visually close species (Bronzed Blackbird vs. Brewer Blackbird), Prompt-CAM needs to look at the eye difference, which is not needed in differentiating visually dissimilar species.
Prompt-CAM can identify and interpret trait manipulation. We conduct a counterfactual-style analysis to investigate whether Prompt-CAM truly relies on the identified traits for making predictions. For instance, to correctly classify the Red-winged Blackbird, Prompt-CAM highlights the red wing patch (the first row of Figure 8), consistent with the field guide provided by the Cornell Lab of Ornithology. When we remove this red spot from the image to resemble a Bronzed Cowbird, Prompt-CAM no longer highlights the original position of the red patch. As such, it does not predict the image as a Red-winged Blackbird but a Bronzed Cowbird (the second row of Figure 8). This shows Prompt-CAM’s sensitivity to trait differences, showcasing its interpretability in fine-grained recognition.
Prompt-CAM uncovers subtle patterns in pair-wise comparison. Following the discussion in subsection 3.4, we illustrate Prompt-CAM’s capability to identify subtle differences between visually similar categories in Figure 9. From the bottom up, the compared classes become more visually similar to the true class, necessitating that Prompt-CAM identifies additional traits for recognition.
@@ 5 Conclusion
We present Prompt Class Attention Map (Prompt-CAM), an effective interpretable method that leverages pre-trained ViTs to identify and localize traits for distinguishing among fine-grained categories. Prompt-CAM is easy to implement and train. Extensive experiments across domains such as animals, plants, and objects demonstrate Prompt-CAM’ superior performance and underscore the potential of repurposing standard models for interpretability.
@@ Appendix A Related Work
Recently, prompting has been introduced in vision transformers (ViTs) to enable efficient adaptation while leveraging the vast capabilities of pre-trained ViTs. Visual Prompt Tuning (VPT) [12] introduces learnable embedding vectors, either in the first transformer layer or across layers, which serve as “prompts” while keeping the backbone frozen. This offers a lightweight and scalable alternative to full fine-tuning, achieving competitive performance on a diverse range of tasks while preserving the pre-trained features.
@@ Appendix F Additional Experiment Results
Ours DINO 97.4 71.9 77.0 87.6 64.7 86.4 99.1 60.8
DINOv2 98.2 73.8 80.0 91.7 70.6 91.9 99.6 62.2

### v2

@@ 3.3 Further Analysis and Discussion
Prompt-CAM on different backbones. Figure 7 illustrates that Prompt-CAM is compatible with different ViT backbones. We show the top three attention maps generated by Prompt-CAM using different ViT backbones on an image of Scott Oriole, highlighting consistent identification of traits for species recognition, irrespective of the backbones. Please see the caption and Suppl. for details.
Prompt-CAM can detect biologically meaningful traits. As shown in Figure 4, Prompt-CAM consistently identifies traits from images of the same species (e.g., the red breast and white belly for Rose-breasted Grosbeak). This is further demonstrated in Figure 1 (d), where we progressively mask attention heads (detailed in subsection 2.3) until the model can no longer generate high-confidence predictions for correctly classifying images of Scott Oriole. The remaining heads 1 and 11 highlight the essential traits, i.e., the black head and yellow belly. Prompt-CAM also enables identifying common traits between species. This is achieved by visualizing the image of one class (e.g., Scott Oriole) using other classes’ prompts (e.g., Brewer Blackbird or Baltimore Oriole). As shown in Figure 1 (c), Brewer Blackbird shares the head and neck color with Scott Oriole. These results demonstrate Prompt-CAM’s ability to recognize species in a biologically meaningful way.
@@
Figure 8: Trait manipulation. The top row shows attention maps for a correctly classified “Red-winged Blackbird” image. In the second row, the red spot on the bird’s wings was removed, and Prompt-CAM subsequently classified it as a “Boat-tailed Grackle,” as depicted in the reference column.
Prompt-CAM can identify and interpret trait manipulation.
We conduct a counterfactual-style analysis to investigate whether Prompt-CAM truly relies on the identified traits for making predictions. For instance, to correctly classify the Red-winged Blackbird, it highlights the red-wing patch (the first row of Figure 8), consistent with the field guide provided by the Cornell Lab of Ornithology. When we remove this red spot from the image to resemble a Boat-tailed Grackle, the model no longer highlights the original position of the red patch. As such, it does not predict the image as a Red-winged Blackbird but a Boat-tailed Grackle (the second row of Figure 8). This shows Prompt-CAM’s sensitivity to trait differences, showcasing its interpretability in fine-grained recognition.
@@ 4 Conclusion
We present Prompt Class Attention Map (Prompt-CAM), a simple yet effective interpretable approach that leverages pre-trained ViTs to identify and localize discriminative traits for fine-grained classification. Prompt-CAM is easy to implement and train. Extensive empirical studies highlight both the strong performance of Prompt-CAM and the promise of repurposing standard models for interpretability.
@@ Appendix A Related Work
Recently, prompting has been introduced in vision transformers (ViTs) to enable efficient adaptation while leveraging the vast capabilities of pre-trained ViTs [53, 12, 42]. Visual Prompt Tuning (VPT) [12] introduces learnable embedding vectors, either in the first transformer layer or across layers, which serve as “prompts” while keeping the backbone frozen. This offers a lightweight and scalable alternative to full fine-tuning, achieving competitive performance on a diverse range of tasks while preserving the pre-trained features.
@@ Appendix F Additional Experiment Results
Ours DINO 98.2 73.2 81.1 91.3 64.7 86.4 99.1 60.8
DINOv2 98.2 74.1 81.3 92.7 70.6 91.9 99.6 62.2